OpenStudy (anonymous):

The density of an object is 450 kg/m^3. What is the mass of the object if the volume is 2.4 m^3? A. 187.5 kg B. 452.4 kg C. 1080 kg D. 6220.8 kg

OpenStudy (phi):

any ideas ? you have something that weighs 450 kg for every m^3 and it has 2.4 m^3

OpenStudy (anonymous):

so, 450*2.4?

OpenStudy (phi):

yes. notice if you look at just the "units" \[ \frac{kg}{m^3} \cdot m^3\] then, like in algebra, we can cancel to get \[ \frac{kg}{\cancel{m^3}} \cdot \cancel{m^3}=kg\] and we get kg... just a weight... which is what we want.

OpenStudy (phi):

in other words, if we divided (for example), the units would not come out correctly... that is a clue to knowing if you should multiply or divide.

OpenStudy (anonymous):

okie dokie. so it would be 450*2.4 = 1,080 which is C

OpenStudy (phi):

yes. and when doing these problems, you should also show the units. It will help you get the correct answer.

OpenStudy (phi):

\[ 450 \frac{kg}{\cancel{m^3}} \cdot 2.4 \cancel{m^3}=1080\ kg\]